Use the technique of completing the square to transform the quadratic equation below into the form (x + c)2 = a.
4x2 + 16x + 12 = 0

Answer by flame8855(424)

4x2 + 16x + 12 = 0
4x^2 + 16x + 16 +12-16=0
(2x+4)^2= 4